OpenStudy (anonymous):

find an equation of the line through 4,-1 and parallel to x+y= 8. use y=mx+b

OpenStudy (anonymous):

first you need to turn x+y=8 into mx+b format, you do this by subtracting x from both sides, that will tell you that the slope of the new line will be 1. Then you use this formula :(y-y1)=m(x-x1) and put in the coordinates to find the answer.

OpenStudy (anonymous):

Can yous how me how

OpenStudy (anonymous):

ok